1. The most recent financial statements for Horn, Inc. are show below (assume no income taxes.) Assets and costs are proportional to sales. Debt and equity are not. No dividends are paid. Next year’s sales are projected to be $8,968. What is the external financing needed (EFN)?

Income Statement    Balance Sheet
Sales    $7,600    Assets    $21,700 Debt $9,100
Costs    $5,180    Equity $12,600
Net Income $2,420 Total    $21,700    Total    $21,700
Answer #1

Growth rate in sales=(8968-7600)/7600=18%

Sales 8968
Costs(5180*1.18) 6112.4
Net income 2855.6

Total assets would be=21700*1.18=25606

Total equity would be=12600+Net income

=12600+2855.6=15455.6

Total assets=Total liabilities+Total equity

Hence external financing needed=25606-(15455.6+9100)

=$1050.4